Chris is planning for her son's college education to begin five years from today. She estimates the yearly tuition, books, and living expenses to be $5,000 per year for a four-year degree. How much must Chris deposit today, at an interest rate of 8 percent, for her son to be able to withdraw $5,000 per year for four years of college?

A. $11,277
B. $13,620
C. $20,000
D. $39,520
1
Expert's answer
2011-10-18T10:46:42-0400
S = P (1 %)^n, 20000 = P 1.08^5, P = 20000/1.08^5 = 13620.
That's why the
correct answer is B: 13620